OpenStudy (anonymous):

identify the slope and the y-intercept of the equation. y =-3/2 x + 4 –3x + 2y = 8

OpenStudy (anonymous):

slope of y =-3/2 x + 4 is -3/2 put x=0, y=4 so, there y-intercept is 4 for –3x + 2y = 8 2y=3x+8 y=(3/2)x+4 so the slope is 3/2 put x=0, y=4 the y-intercept is 4

OpenStudy (anonymous):

Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.